OpenStudy (anonymous):

In comparing two positive numbers expressed in scientific notation can one assume that the number with the greater exponent is automatically the greater number

OpenStudy (anonymous):

what about \[2^3~ and~ 3^2\]

OpenStudy (anonymous):

@surjithayer \( 2^3 and\space 3^2 \) is not scientific. Scientific has a 10 to the power of something. So \(2.23 \times 10^2 \) is Scientific notation.

OpenStudy (anonymous):

o kay ,i forget this statement

OpenStudy (anonymous):

So if you compare Scientific notation you will be comparing numbers like \(\huge 2.3 \times 10^2 \) and \(\huge 2.3 \times 10^5 \) which is greater or lesser??

OpenStudy (anonymous):

sorry i have not noticed the word scientific

OpenStudy (anonymous):

First you compare the exponents. The one with the greater exponent will be larger than the lesser exponent one. if they have the same exponent, you look at the number and compare the number

OpenStudy (anonymous):

take the case of approx. two extreme positins \[1.1*10^2=110\] \[9.9*10^1=99\] so it concludes that one with greater exponent is greater.